OpenStudy (anonymous):

Which expression shows log3(3x^4) in expanded form? log3^3 + 4log3^x 4log3^3x 4log3^3 + 4log3x 3log3^3 + 4log3x

Nnesha (nnesha):

quotient rule\[\large\rm log_b x - \log_b y = \log_b \frac{ x }{ y}\] to condense you can change subtraction to division product rule \[\large\rm log_b x + \log_b y = \log_b( x \times y )\] addition ----> multiplication power rule \[\large\rm log_b x^y = y \log_b x\]

OpenStudy (anonymous):

I don't really understand this

Nnesha (nnesha):

you just need to apply those properties

OpenStudy (anonymous):

would it be the first one?

Nnesha (nnesha):

\[\huge\rm \log_3(3 · x^4) \] 3 times x^4 multiplication so which property would u apply ?

Nnesha (nnesha):

how did you get first first one ?

Nnesha (nnesha):

which property did u use ? :=)